Reports 1-1 of 1 Clear search Modify search
AdV-ALS (Installation)
gosselin, derossi, sposito, montanari - 16:21 Wednesday 09 June 2021 (52090) Print this report
New fiber for IR pick off

This morning we plugged the new fiber for the IR pick off (PSL towards ALS/SQZ/phase camera) that we installed last week with Beatrice and Gianmatteo. It is 10 meters long and passes in the canal duct next to the ceiling.
This is a fiber from oz optics having a metallic protection (see picture). We will check with the next lock if it helps in reducing the noise that can be seen in the green error signals CARM and DARM.

We realigned the beam into the fiber to maximize the coupling and decreased the power at its input to get back to the 134 mW that we had at the input of the splitter box.

To be noted that during the intervention we accidentaly interupt the beam seeding the fiber amplifier and this last one went into security. We switched it on again and closed all the loops of the injection.

Images attached to this report
Comments to this report:
derossi, gosselin - 17:41 Thursday 10 June 2021 (52110) Print this report

Giving a look at the data it seems that there is an improvement between 10 and 100 Hz on CARM and DARM error signals. To be analyzed more in detail.

Images attached to this comment
Search Help
×

Warning

×